Not sure on how to go about this problem.
4(Exponet 2x)=9(exponet x-1). solve for x

Answer by stanbon(57347)

4(Exponet 2x)=9(exponet x-1). solve for x
4^(2x) = 9^(x-1)
Take the log of both sides to get:
(2x)(log4) = (x-1)(log9)
2xlog4=xlog9-log9
2xlog4-xlog9=-log9
x(2log4-log9)=-log9
x(0.24987747) = -0.95424251
x=-3.81884173...
Cheers,
Stan H.
Comment: There are many other ways to solve this
problem. Your instructor may have used a different
method.